OpenStudy (anonymous):

PLEASE HELP!! If x =3 and y = -2 , then y4 - 4x = ?

OpenStudy (anonymous):

16-12 = 4

OpenStudy (angela210793):

(-2)^4-4*3=16-12=4

OpenStudy (katrinakaif):

-2^4 - 4x 3 = 16 - 12 = 4

OpenStudy (anonymous):

-2x4 + 3x4 -8 + 12 4

OpenStudy (anonymous):

no cuz its (-2)^4 no (-2)*4

OpenStudy (anonymous):

thank you everyone !! :))

OpenStudy (anonymous):

hey i tried lol